OpenStudy (anonymous):

Hello! :-) What is the equation in standard form of the line which passes through (4, −2) and has a slope of −3? Answer Options: a. x + 3y = 10 b. x + 3y = −10 c. 3x + y = 10 d. 3x + y = −10 Thanks in advance! :-)

563blackghost (563blackghost):

We first need to put the pair and the slope into y=mx+b....

OpenStudy (misty1212):

Hello!

OpenStudy (anonymous):

Answer is: c. 3x + y = 10.

OpenStudy (misty1212):

ok

OpenStudy (misty1212):

would you like to know how to get it?

OpenStudy (anonymous):

the general equation of the slope is given by \[y -y _{o}=m(x - x _{o})\] where \[y _{o} = -2\] and \[x _{o}= 4\] and slope = m = -3 by solving the equation we get \[3x +y =10\]

OpenStudy (anonymous):

the line*** not the slope